OpenStudy (anonymous):

If 3:sqrt(t^2)+2sqrt(t^3) find f'(t)

OpenStudy (anonymous):

\[f(x)=\sqrt{t^2}+2\sqrt{t^3}\]?

OpenStudy (anonymous):

the first square root is ^3

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

\[\sqrt[3]{t^2}+2\sqrt{t^3}\]

OpenStudy (anonymous):

\[\sqrt[3]{t^2}+\sqrt{t^3}\] maybe

OpenStudy (anonymous):

got it rewrite as \[t^{\frac{2}{3}}+2t^{\frac{3}{2}}\] and use the almighty power rule (once for each term)

OpenStudy (anonymous):

2/3t^(-1/3)+3t^(1/2)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw might want to write as \[\frac{2}{3\sqrt[3]{t}}+3\sqrt{t}\]

OpenStudy (anonymous):

oh okay thanks.